### v1

# thanks: This work is based on the author’s M.Sc. dissertation (Ref.[1]), but further correcting and expanding upon it

Abstract
Abstract
We derive the closed-form one-loop Euler–Heisenberg effective actions for Dirac fermions coupled simultaneously to classical electromagnetic vector and massive pseudo-vector backgrounds within a controlled quasi-static approximation. Through complete diagonalization of the functional Hessian, we systematically delineate the parameter space into distinct sectors characterized by stability properties and spectral structure. We identify subspaces that encompass and extend results from previous studies into a broader class, admitting propagating axial fields as physically viable regimes; strikingly, we note a sector presenting chirality-asymmetric instability. This addresses long-standing questions regarding the well-defined nature, diagonalizability, and stability of the model. From the effective action, we derive novel nonperturbative pair-production rates for simultaneously propagating electromagnetic and axial vector backgrounds; remarkably, we find pronounced vacuum stabilization compared to previous results. Furthermore, we demonstrate that this framework allows for a unified derivation of the chiral anomaly structures in the general case and show that the electromagnetic coupling induces instanton-like configurations for the axial field, even when it is not a fundamental gauge field. As a proof-of-concept, we analyze a cosmological toy model of baryogenesis driven by an axial vector, providing numerical estimates that support the viability of this hypothesis. Additionally, we outline qualitative predictions for Weyl/Dirac semi-metals and briefly discuss potential applications in related phenomena, such as the Strong-CP problem.
Euler–Heisenberg effective action; Schwinger effect; axial vector background; Weyl/Dirac materials; Hessian diagonalization
pacs:
12.20.-m, 11.30.Rd, 11.10.Gh, 72.80.Vp
I Introduction
The nonperturbative response of the quantum vacuum to strong classical fields is a cornerstone of quantum electrodynamics (QED). Integrating out charged fermions yields the Euler–Heisenberg (EH) effective action [2], which encodes low-energy ($\leq mc^{2}$) nonlinear photon interactions and underpins quantitative predictions for Schwinger pair production [3], vacuum birefringence, photon-photon scattering [4], and related phenomena. For homogeneous electromagnetic backgrounds, closed-form EH expressions are routinely derived via proper-time [5] or numerical worldline [6] methods, establishing the standard framework for field-induced particle production and nonlinear optics [7, *exactplanewave, *exactselfdual1, *exactselfdual2, *exactsauter, *exactplanewave2]. In recent decades, this formalism has been extended to include fermions coupled simultaneously to a vector potential $A_{\mu}$ and a parity-violating axial-vector field $S_{\mu}$—a configuration that arises in three distinct physical arenas: (i) gravitational theories with torsion, where the pseudo-trace of the contorsion tensor acts as $S_{\mu}$ and can seed baryogenesis [13, *stringtorsion, 15, *ShapiroBelyaevTorsion, 17, *KalbRamond, *2Bornot2B]; (ii) high-energy extensions of the Standard Model, where axial fields may catalyze magnetic monopole pair production [20, *monopole]; and (iii) Weyl/Dirac semimetals, where momentum-space node separations modulated by strain or magnetic textures generate emergent axial couplings [22, *torsionweylmetal, *Topologicalmaterials]. Despite their disparate origins, these contexts share a common field-theoretic core, motivating a unified nonperturbative treatment capable of bridging cosmological dynamics and condensed-matter experiments.
Pioneering work by Maroto [25, 26] first revealed that constant axial backgrounds in quasi-static configurations ($\partial_{\mu}S^{\mu}=0$) can suppress Schwinger pair production—a stark departure from pure QED. Subsequent studies generalized the EH action to vector-axial systems, but often yielded operator-valued expressions requiring further contractions, limiting practical utility [27]. Recent advances identified special field configurations—e.g., constant, light-like, or parallel $S_{\mu}$—that admit exact diagonalization and exhibit either enhanced or suppressed pair creation [26, 28, *Copinger1, 30]. Concurrently, condensed-matter theory has clarified how strain and Floquet driving emulate axial fields in topological semimetals, opening pathways to test Lorentz violation and nonlinear axial responses in lab [31, 32, *Diractorsion, 34, *Lorentzviolationsemimetals2]. Yet a systematic classification of the full parameter space—delineating when the model is stable, Hermitian, and diagonalizable—has remained absent. The core obstacle lies in the functional Hessian: axial-vector interference can render it non-Hermitian or defective, signaling vacuum instability. Moreover, consistency conditions from torsion-motivated UV completions (e.g., Planck-scale mass, ghost-freedom [36]) further constrain viable regimes, demanding a rigorous sector-by-sector analysis of physical admissibility.
In this work, we present a complete one-loop analysis of models containing fermions simultaneously coupled to a vector and a possibly massive pseudo-vector fields in configurations with constant field invariants (quasi-static/adiabatic settings), which we shall call electroaxial theories in what follows. Our work builds upon and extends the foundations laid before through the Following contributions: (i) The Complete Hessian diagonalization and systematic parameter-space classification of well-defined and stable subspaces; (ii) The derivation of Exact closed-form effective actions within the viable sectors even for propagating axial-vector field; (iii) The unified derivation of anomaly structures and demonstration of constraints that forces the pseudo-vector to inherit topological structures from the gauge vector field; (iv) The extraction of novel Pair-production rates for propagating axial-vector backgrounds and (v) The Phenomenological applications with numerical estimates for relic field strength from a proof-of-concept toy-model for baryogenesis and qualitative assessments for condensed matter and other possible applications. Remarkably, our results indicate a pronounced vacuum-stabilization when the pseudo-vector is present and estimate bounds for the viability of hypotheses in which baryogenesis is driven by an axial-vector.
The remainder of this paper is organized as follows. Section II reviews the theoretical framework: the electroaxial action and proper-time representation. Section III presents the complete Hessian diagonalization and parameter-space classification, establishing stability criteria for each sector. Section IV derives the exact one-loop effective action in closed form through zeta-function regularization, with detailed consistency checks against perturbative expansions and known limits. Section V analyzes the anomaly structure and topological constraints on axial vector configurations. Section VI extracts vacuum-transition-probability measures and pair-production rates, including the new results for propagating axial fields and explores phenomenological implications: cosmological baryogenesis (Sec. VI.2) and further applications (Sec. VI.3) with qualitative predictions for condensed matter and brief remarks on connections to the strong CP problem, dark energy, and current experimental searches for magnetic monopoles. We conclude in Section VII.
Throughout, we employ natural units ($\hbar=c=1$), the Minkowski metric signature (+,,,)(+,-,-,-), and adopt conventions consistent with Refs. [37, *QFTQGdoprof] for spinor algebra. Numerical estimates use $M_{\mathrm{Planck}}=1.22\times 10^{19}\,\mathrm{GeV}$ and Planck 2018 cosmological parameters [39].
II The Electroaxial Theory
The strategy adopted throughout closely follows the approach established by Maroto [26], which is itself an extension of Schwinger’s seminal treatment of the pure QED case [3]—a procedure systematically reviewed in Refs. [40, 7]. We assume a familiarity with functional methods at the level of standard quantum field theory texts [41, 38] while providing sufficient detail to make the calculations self-contained and establish notation conventions consistent with the broader literature on background-field effective actions.
We analyze a model-dependent theory of Dirac fermions $\psi$ coupled concurrently to an electromagnetic vector potential $A_{\mu}$ and a (possibly massive) pseudo-vector field $S_{\mu}$. The classical action in Minkowski spacetime is expressed as
$\mathrm{S}\left[A,S,\,\overline{\!{\psi}},\psi\right]=\mathrm{S}_{0}\left[A,S\right]+\mathrm{S}_{\text{Dirac}}$ (1)
Here, $\mathrm{S}_{0}\left[A,S\right]$ pertains only to the vector fields, incorporating the standard Maxwell action
$\mathrm{S}_{\text{Maxwell}}=-\frac{1}{4}\int d^{4}xF_{\mu\nu}F^{\mu\nu}$ (2)
along with the free axial-vector action [15, *ShapiroBelyaevTorsion]
$\mathrm{S}_{\text{Proca}}=\int d^{4}x\left[-\frac{1}{4}S_{\mu\nu}S^{\mu\nu}+\frac{1}{2}M_{S}^{2}S_{\mu}S^{\mu}\right]$ (3)
In these expressions, $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ and $S_{\mu\nu}=\partial_{\mu}S_{\nu}-\partial_{\nu}S_{\mu}$ are the field-strength tensors corresponding to the vector and pseudo-vector sectors, respectively, while $M_{S}$ denotes the mass of the pseudo-vector field. Based on the model of choice, classical interaction terms (e.g., $A_{\mu}S^{\mu}$) can also be included in the analysis.
The Dirac action for the fermions is defined as
[…]
$V_{\mu}^{(R/L)}=eA_{\mu}\pm\eta S_{\mu}$ (5)
rather than to a common gauge field. This allows for the convenient notation $D_{\mu}=\partial_{\mu}-iV^{(R)}_{\mu}\mathrm{P}_{R}-iV^{(L)}_{\mu}\mathrm{P}_{L}$ for the covariant derivative.
The effective action can then be constructed by integrating out the fermions within the generating functional:
$Z\left[A,S\right]=\frac{\int\mathcal{D}\,\overline{\!{\psi}}\mathcal{D}\psi e^{i\mathrm{S}\left[A,S,\,\overline{\!{\psi}},\psi\right]}}{\int\mathcal{D}\,\overline{\!{\psi}}\mathcal{D}\psi e^{i\mathrm{S}\left[A=S=0\right]}}=e^{iW\left[A,S\right]}$ (6)
This yields the effective action
[…]
The one-loop quantum corrections are encapsulated in the last term, which forms the focal point of our analysis.
For simplicity, we assume that the background fields are smooth and confined to a compact spatial domain, with appropriate boundary conditions, within which the quasi-static approximation holds [42]. This assumption is standard for EH-type effective actions [3, 7] as it enables the application of translation-operator manipulations described in Sec. IV while effectively capturing the physics of slowly varying fields. Generalizations to time-dependent backgrounds extend beyond one-loop exactness and typically require resummation techniques [43] or numerical worldline methods [6]. Under this quasi-static approximation, the extended Dirac operator exhibits elliptic properties and possesses a discrete spectrum, thereby ensuring it is trace-class [5, *HeatKernel2]. Consequently, the traces of $\not${D} and its transpose are equivalent, such that
[…]
The transpose of the Dirac operator can be derived using the charge conjugation matrix defined as $C=\gamma^{2}\gamma^{0}$. This matrix exhibits the properties:
$C\gamma^{\mu}C^{-1}=-\left(\gamma^{\mu}\right)^{\mathrm{T}}$ and $C\gamma^{5}C^{-1}=\left(\gamma^{5}\right)^{\mathrm{T}}=\gamma^{5}$ (9)
These properties yield:
[…]
Adding the term on the left-hand side, we find that this correction can be represented as:
[…]
In this equation, we denote the quantum corrections as $\overline{\Gamma}\left[A,S\right]$. Direct substitution gives:
[…]
Thus, we arrive at:
[…]
Here, we have utilized Eq. (8) and substituted Eq. (13), with both numerator and denominator multiplied by 1-1 for convenience.
Next, we map this result into operator formalism as follows, where we adopt the hat notation for operators temporarily. Let $\hat{x}^{\mu}$ and $\hat{p}_{\mu}$ be the position and momentum operators respectively which obey the commutation relations $\left[\hat{x}^{\mu},\hat{p}_{\nu}\right]=i\delta^{\mu}_{\nu}$, each possessing a complete eigenbasis $\hat{x}^{\mu}\left|x\right>=x^{\mu}\left|x\right$> and $\hat{p}_{\mu}\left|p\right>=p_{\mu}\left|p\right$> satisfying
Next, we transition to an operator formalism, employing hat notation for operators temporarily. Let $\hat{x}^{\mu}$ and $\hat{p}_{\mu}$ be the position and momentum operators, respectively, which satisfy the commutation relations $\left[\hat{x}^{\mu},\hat{p}_{\nu}\right]=i\delta^{\mu}_{\nu}$, each possessing a complete eigenbasis $\hat{x}^{\mu}\left|x\right>=x^{\mu}\left|x\right$> and $\hat{p}_{\mu}\left|p\right>=p_{\mu}\left|p\right$> satisfying
$\left<p\vphantom{x}\right|\left.x\vphantom{p}\right>=\frac{1}{\left(2\pi\right)^{2}}e^{-ip\cdot x}$ (15)
Therefore, we have:
$\left<x\right|\hat{p}_{\mu}\left|\Psi\right>=-i\partial_{\mu}\left<x\vphantom{\Psi}\right|\left.\Psi\vphantom{x}\right$> (16)
for any state $\left|\Psi\right$>. We denote
$i\hat{D}_{\mu}\left<x\vphantom{\Psi}\right|\left.\Psi\vphantom{x}\right>\to\left|\Psi\right>=\left<x\right|\left(\hat{\Pi}_{\mu}+\eta\hat{S}_{\mu}\gamma^{5}\right)\left|\Psi\right$> (17)
where $\hat{\Pi}_{\mu}=\hat{p}_{\mu}-e\hat{A}_{\mu}$, and we will drop the hat notation for the remainder of this paper. By employing the identities [45]:
$\gamma^{5}\sigma^{\mu\nu}=-\frac{i}{2}\varepsilon^{\mu\nu\alpha\beta}\sigma_{\alpha\beta}$ and $\gamma^{5}\gamma^{\mu}=-\frac{1}{6}\varepsilon^{\mu\nu\alpha\beta}\gamma_{\nu}\sigma_{\alpha\beta}$ (18)
we simplify Eq. (14) to
$\overline{\Gamma}[A,S]=-\frac{i}{2}\mathrm{Tr}\ln\left[\frac{\left(\Pi_{\mu}+\eta S_{\mu}\gamma^{5}\right)^{2}-\frac{1}{2}\bigg(V_{\mu\nu}+\frac{2\eta m}{3}\varepsilon_{\alpha\beta\mu\nu}S^{\alpha}\gamma^{\beta}\bigg)\sigma^{\mu\nu}-m^{2}}{p^{2}-m^{2}}\right],$ (19)
where we define $V_{\mu\nu}=eF_{\mu\nu}-i\eta\tilde{S}_{\mu\nu}$ and $\tilde{S}_{\mu\nu}=\frac{1}{2}\varepsilon_{\alpha\beta\mu\nu}S^{\alpha\beta}$.
Now, we proceed to employ Fock–Schwinger’s proper-time representation [3, 46] for every $\tau>0$, provided that analytic continuation exists:
$\overline{\Gamma}\left[A,S\right]=\frac{i}{2}\int_{0}^{\infty}\frac{d\tau}{\tau}e^{-i\tau m^{2}}\mathrm{Tr}\left[e^{-i\tau\mathcal{H}}-e^{-i\tau p^{2}}\right]$ (20)
where $\mathcal{H}$ denotes the quadratic Hessian functional:
$\mathcal{H}=\left(\Pi_{\mu}+\eta S_{\mu}\gamma^{5}\right)^{2}-\frac{1}{2}\bigg(V_{\mu\nu}+\frac{2\eta m}{3}\varepsilon_{\alpha\beta\mu\nu}S^{\alpha}\gamma^{\beta}\bigg)\sigma^{\mu\nu}$ (21)
This representation brings forth a long and well discussed [3, 41, 38] analogy between these exponentials in Eq. (20) and time-evolution operators such as $e^{-i\tau\mathcal{H}}\left|x\right>=\left|x;\tau\right$>. In this context, $\tau$ acts as a form of “proper time,” while the remaining terms function akin to a “Hamiltonian” that adheres to its own Schrödinger-like functional equations, expressed as follows:
$i\partial_{\tau}\left<y;0\vphantom{x;\tau}\right|\left.x;\tau\vphantom{y;0}\right>=\left<y;0\right|\mathcal{H}\left|x;\tau\right$> (22)
Here, $\partial_{\tau}$ denotes the partial derivative with respect to $\tau$.
Moreover, the correlation between the last term in Eq. (20) and the coincidence limit of the free particle’s propagator facilitates the application of standard integration techniques within momentum representation [47]. This yields:
$\mathrm{Tr}\ e^{-i\tau p^{2}}=\int d^{4}x\left<x\right|e^{-i\tau p^{2}}\left|x\right>\mathrm{tr}\mathbf{1}=\lim_{y\to x}\frac{4}{(2\pi)^{4}}\int d^{4}x\ d^{4}p\left<y;0\vphantom{p}\right|\left.p\vphantom{y;0}\right>\left<p\vphantom{x;\tau}\right|\left.x;\tau\vphantom{p}\right>=\int d^{4}x\ \frac{i}{(2\pi\tau)^{2}}$ (23)
In this equation, $\mathrm{tr}\mathbf{1}=4$ represents the trace in Dirac space. In contrast, the first term in Eq. (20) presents complexities due to its Hessian, which necessitates a detailed analysis that will be conducted in the following section.
III Hessian Diagonalization and Sector Classification
In this section, we present the diagonalization procedure and the taxonomy of sectors. We begin by decomposing the first term of the Hessian as follows:
$(\Pi+\eta S\gamma^{5})^{2}=\left(\Pi^{2}+\eta^{2}S^{2}\right)\mathbf{1}+2\eta\Pi_{\mu}S^{\mu}\gamma^{5}$ (24)
Here, $S^{2}=S_{\mu}S^{\mu}$, and it is noteworthy that any operator product will henceforth be interpreted as anti-commutators; for instance,
$2\Pi_{\mu}S^{\mu}\equiv\left\{\Pi_{\mu},S^{\mu}\right\}=\Pi_{\mu}S^{\mu}+S^{\mu}\Pi_{\mu}=i\partial_{\mu}\left(S^{\mu}\right)+2eA_{\mu}S^{\mu}$ (25)
We can recast (21) as $\mathcal{H}=\left(\Pi^{2}+\eta^{2}S^{2}\right)\mathbf{1}+\mathcal{H}_{\mathrm{R}}$. The remaining term is defined as:
$\mathcal{H}_{\mathrm{R}}=2\eta\Pi_{\mu}S^{\mu}\gamma^{5}-\frac{1}{2}\bigg(V_{\mu\nu}+\frac{2\eta m}{3}\varepsilon_{\alpha\beta\mu\nu}S^{\alpha}\gamma^{\beta}\bigg)\sigma^{\mu\nu}$ (26)
This term features a characteristic equation [*[ObtainedbymeansofaPythoncodewrittenbytheauthorandsolvedusingWolfram'sMathematica.Check][forfurtherdetails.]diagmine],
$\lambda^{4}+4\mathcal{A}\lambda^{2}+16i\mathcal{B}\lambda+4\mathcal{C}=0$ (27)
The definitions of its elements can be found in the Appendix. The solutions take the form:
$\lambda^{(\pm)}_{\pm}=\left(\pm\right)\frac{\sqrt{6}}{6}\left[\sqrt{\mathcal{D}_{-}}\pm i\sqrt{\mathcal{D}_{+}\left(\pm\right)\frac{24i\mathcal{B}}{\sqrt{\mathcal{D}_{-}}}}\right]$ (28)
where the notations $(\pm)$ and $\pm$ represent independent signs, with $\mathcal{D}_{\pm}$ [Eq. 103] being complicated functions of:
[…]
with the usual fundamental invariants of the electromagnetic field denoted as $\mathcal{F}_{A}=\frac{1}{4}F_{\mu\nu}F^{\mu\nu}$ and $\mathcal{G}_{A}=-\frac{1}{4}\tilde{F}^{\mu\nu}F_{\mu\nu}$ while the axial-vector analogues are specified as $\mathcal{F}_{S}=\frac{1}{4}S_{\mu\nu}S^{\mu\nu}$ and $\mathcal{G}_{S}=-\frac{1}{4}\tilde{S}^{\mu\nu}S_{\mu\nu}$.
From Eqs. (28) and (103), we observe that there exist certain points in the parameter space at which the Hessian ceases to be Hermitian. Specifically, this occurs when $\mathcal{D}_{-}\to 0$ or $\infty$, indicating configurations where (i) $\Delta\to 0$ [Eq. 104], (ii) $\Delta\to\infty$, (iii) $\mathcal{A}\to\infty$, and (iv) $\mathcal{B}\to\infty$ are unstable, suggesting potential symmetry breaking or phase transitions [49, *degeneracies, *PhaseTransitions]. Additionally, there are regions where the imaginary part of the eigenvalues may be relevant or even dominant. While this may initially seem unphysical for effective actions derived from fundamental theories (such as Quantum Gravity and Extensions of the Standard Model), it could prove beneficial for deriving particle creation rates within these theories, or be applicable in other domains, such as Open Systems [52] and PT-Symmetric models [53], which are useful in Quantum Optics [54], Nuclear Physics [55, *ComplexHamiltonianApplication], and beyond.
The parameter space can be systematically categorized into various types based on the limiting regions where the system displays distinct properties:
Extended-Quantum-Electrodynamics-type (EQED):
In this regime, the conditions $\Pi_{\mu}S^{\mu}=\widetilde{V}^{\mu\nu}S_{\nu}=0$ and $S^{2}\simeq 0$ hold, leading to the simplification of the Hessian to $\mathcal{H}_{\mathrm{R}}=-\frac{1}{2}V_{\mu\nu}\sigma^{\mu\nu}$. This yields expressions that are analogous to those found in QED, where the quantity $V^{\mu\nu}$ serves as an extended version of the Faraday tensor $F_{\mu\nu}$. The eigenvalues associated with this system can be expressed as follows:
$\lambda^{(\pm)}_{\pm}=(\pm)\sqrt{2\mathcal{V}\pm 2i\mathcal{W}}$ (30)
Mixed-Transverse-type (MT):
In the MT case, the fields are constrained by the condition $\Pi_{\mu}S^{\mu}=0$, indicating that they are transverse and orthogonal, specifically $A_{\mu}S^{\mu}=\partial_{\mu}\left(S^{\mu}\right)=0$. Under these constraints, the eigenvalues take the following form:
$\lambda^{(\pm)}_{\pm}=\left(\pm\right)\sqrt{2\mathcal{V}-4m^{2}\eta^{2}S^{2}\pm 2\sqrt{4m^{2}\eta^{2}\left(\widetilde{V}^{\mu\nu}S_{\nu}\right)^{2}-\mathcal{W}^{2}}}$ (31)
The system remains well-behaved as long as the condition $\mathcal{V}\geq 2m^{2}\eta^{2}S^{2}\pm\Re\sqrt{4m^{2}\eta^{2}\left(\widetilde{V}^{\mu\nu}S_{\nu}\right)^{2}-\mathcal{W}^{2}}$ is satisfied. It is noteworthy that the EQED-type is encompassed within this broader subspace.
Non-anomalous-type (NonAnom):
In this region, where $\mathcal{W}=0$, we find the eigenvalues given by:
$\lambda^{(\pm)}_{\pm}=\left(\pm\right)\sqrt{2\mathcal{V}+4\eta^{2}\left[\left(\Pi_{\mu}S^{\mu}\right)^{2}-m^{2}S^{2}\right]\pm 4\eta\sqrt{2\left(\Pi_{\mu}S^{\mu}\right)^{2}\mathcal{V}+m^{2}\left(\widetilde{V}^{\mu\nu}S_{\nu}\right)^{2}}}$ (32)
This regime corresponds to disconnected vector and pseudo-vector sectors, which propagate in configurations for which any local CP violation arising from one field is counterbalanced by an opposite CP violation from the other field. This dynamic results in an (almost) preserved CP symmetry; for instance, a circularly polarized light ray is associated with an “axial-light ray” from $S_{\mu}$ in the opposite polarization [*[FurtherdiscussionofthiskindofbehaviorcanbefoundatsectionsD.3andD.4of][.]Hehlelectro]. When $S_{\mu}$ is massive, CP violations are confined to its characteristic scale ($\sim 1/M_{S}$), thus allowing CP conservation to be inherited by the freely propagating $A_{\mu}$ at larger distances.
However, there are important considerations: the UV calibration of the theory must be approached with caution to avoid deeper issues such as ghosts [36]. Generally, the underlying model cannot be fundamental, and the mass of the axial vector must be significantly larger than that of the fermion to prevent instability in the IR sector. For a very massive axial-vector, given the equations (1), (3), and (4), along with the quasi-static assumption at low energies ($\partial_{\mu}S^{\mu\nu}\approx 0$), we then have:
$\frac{\delta\mathrm{S}\left[A,S\right]}{\delta S_{\mu}}\approx S^{\mu}+\frac{1}{M^{2}_{S}}j_{5}^{\mu}=0$ (33)
where $j_{5}^{\mu}=\left<\,\overline{\!{\psi}}\gamma^{\mu}\gamma^{5}\psi\right$>. This leads us to conclude that:
$\left|\Pi_{\mu}S^{\mu}\right|\sim\frac{m^{2}}{M^{2}_{S}}\approx 0$ (34)
From this, we can further simplify the eigenvalues to:
$\lambda^{(\pm)}_{\pm}\approx\left(\pm\right)\sqrt{2\mathcal{V}-4\eta^{2}m^{2}S^{2}\pm 4m\eta\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|}$ (35)
These eigenvalues remain stable under the condition $\mathcal{V}\geq 2\eta m\left(\eta mS^{2}+\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|\right)$ and are amenable to analysis since they avoid the complication of handling $\left(\Pi_{\mu}S^{\mu}\right)$ within a square root.
Teleparallel Rainich–Misner–Wheeler-type (TRMW):
This subspace is contained within the previous one, but it introduces the additional constraint of $\mathcal{V}=0$. In a teleparallel or Einstein–Cartan setting, where the axial vector is derived from a torsion tensor ($S_{\mu}=\varepsilon_{\mu\nu\alpha\beta}T^{\mu\nu\alpha\beta}$ [15, *ShapiroBelyaevTorsion]), this constraint realizes a Rainich-style [58] identification of an electromagnetic-like structure with torsion. Notably, this constraint also necessitates the self-duality of the fields, expressed as $\widetilde{V}^{\mu\nu}=\pm iV^{\mu\nu}$.
Apart from this, Eq. (32) simplifies to
$\lambda^{(\pm)}_{\pm}=\left(\pm\right)2\sqrt{\eta^{2}\left(\Pi_{\mu}S^{\mu}\right)^{2}-\eta^{2}m^{2}S^{2}\pm\eta m\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|}$ (36)
When the condition $\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|=0$ is satisfied, the Hessian of a system lying within this subspace can be hermitian if and only if $mS\leq\left|\Pi_{\mu}S^{\mu}\right|$. A special simplification arises for a pseudo-vector in light-like configurations ($S^{2}=0$), which can be achieved either (i) classically if it is massless or (ii) off-shell if it is quantized. In such cases, the Hessian takes the form:
$\mathcal{H}=\begin{bmatrix}\left(\Pi_{\mu}+\eta S_{\mu}\right)^{2}\mathbf{1}&0\\ 0&\left(\Pi_{\mu}-\eta S_{\mu}\right)^{2}\mathbf{1}\end{bmatrix}$ (37)
where $\mathbf{1}$ denotes a 2-by-2 identity matrix block. If $\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|\neq 0$ but $M_{S}\gg m$, we find
$\mathcal{H}=\begin{pmatrix}\left[\left(\Pi_{\mu}+\eta S_{\mu}\right)^{2}+2\sqrt{\eta m\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|}\right]\mathbf{1}&0\\ 0&\left[\left(\Pi_{\mu}-\eta S_{\mu}\right)^{2}+2i\sqrt{\eta m\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|}\right]\mathbf{1}\end{pmatrix}$ (38)
which exhibits an unstable vacuum for one chirality and conversely for the other. In time-like configurations of the axial vector, the Hessian exhibits these same signs of instability, even for a very massive axial vector, leading to the expression:
$\lambda^{(\pm)}_{\pm}\approx\left(\pm\right)2\sqrt{\eta m\left(-\eta mS^{2}\pm\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|\right)}$ (39)
This implies that the eigenvalues invariably contain at least two pure imaginary components, provided the condition $\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|\geq\eta mS^{2}$ holds. If attainable, this scenario indicates a chirality-asymmetric stability, where parity is violated in a manner that allows one chirality sector to remain “safe”  [^1] This safety can be put into question if higher loops, interactions, and backreactions are included according to more complicated scenarios. while the opposite sector invariably decays into particles. This interplay results in a chirality-asymmetric pair creation, characterized by a net production of fermions with a preferred chirality. Should any process propel $\eta mS^{2}$ beyond the remaining term, the system will become entirely unstable, precipitating rapid vacuum decay.
Significantly, this sector also demonstrates a kinematical symmetry reduction to the $SIM(2)$ (or $ISO(2)$ if the norm of $S_{\mu}$ is fixed) subgroup of the Lorentz group. This assertion is underscored by the structure of Eq. (37), which bears resemblance to that found in Very Special Relativity [60]. Similar symmetry breaking is also observed in the preceding subspaces, though not with the same level of explicitness and specificity. Consequently, models constructed within this subspace have the potential to exhibit a rich phenomenology, including the intriguing possibility of imparting mass to neutrinos without necessitating a Seesaw mechanism [61], provided that careful considerations are made during the model’s development [*[WhiletheHessianinquestioniswell-behaved, itmaynotencompassalldesiredrepresentations, includingprojectiveones, asdiscussedby][.]Marcondes, *Marcondes2].
Other types:
Apart from the previously discussed subspaces, further stable regions may exist elsewhere in parameter space, contingent upon careful considerations and specific choices informed by the structure of Eq. (28). Given that the combined regions of MT- and NonAnom-types encompass a significant portion of the parameter space, the remaining regions are either special cases within/transitional between these two categories—such as those involving the vanishing of additional invariants or potentially exhibiting complex eigenvalue structures—or intrinsically unstable for any parameter values. Noteworthy regions that fall under the category of exhibiting long eigenvalue expressions are strictly governed by conditions such as $S^{\mu}=0$, $\mathcal{V}=0$, $\left(V^{\mu\nu}S_{\nu}\right)^{2}=0$, or any combinations thereof, without imposing additional constraints on other invariants.
Type $\left(\lambda^{(\pm)}_{\pm}\right)^{2}$ Stability Theory Accessibility
EQED $2\mathcal{V}\pm 2i\mathcal{W}$ $\sim$ QED’s conditions Massless $S_{\mu}$
MT $2\mathcal{V}-4m^{2}\eta^{2}S^{2}\pm 2\sqrt{4m^{2}\eta^{2}\left(\widetilde{V}^{\mu\nu}S_{\nu}\right)^{2}-\mathcal{W}}$
$\mathcal{V}\geq 2m^{2}\eta^{2}S^{2}\pm$
$\pm\Re\sqrt{4m^{2}\eta^{2}\left(\widetilde{V}^{\mu\nu}S_{\nu}\right)^{2}-\mathcal{W}^{2}}$
All theories
NonAnom
$2\mathcal{V}+4\eta^{2}\left[\left(\Pi_{\mu}S^{\mu}\right)^{2}-m^{2}S^{2}\right]\pm$
$\pm 4\eta\sqrt{\left(\Pi_{\mu}S^{\mu}\right)^{2}\mathcal{V}+m^{2}\left(\tilde{V}^{\mu\nu}S_{\nu}\right)}$
$m\ll M_{S}$ plus same as MT Massive $S_{\mu}$
TRMW $4\left[\eta^{2}\left(\Pi_{\mu}S^{\mu}\right)^{2}-\eta^{2}m^{2}S^{2}\pm\eta m\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|\right]$
$\eta\left(\Pi_{\mu}S^{\mu}\right)^{2}\geq\eta m^{2}S^{2}\pm m\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|$
Possibly Chirality-Asymmetric-Unstable
All theories
(Better for massless $S_{\mu}$)
Others $\frac{1}{6}\left[\sqrt{\mathcal{D}_{-}}\pm i\sqrt{\mathcal{D}_{+}\left(\pm\right)\frac{24i\mathcal{B}}{\sqrt{\mathcal{D}_{-}}}}\right]^{2}$ Needs careful handling All theories
Table 1: Summary of the sector taxonomy presented here.
We conclude this section by noting that the MT-type represents a consistent subspace for both massive and massless axial vectors, provided the inequality
$\mathcal{V}\geq 2\eta^{2}m^{2}\eta mS^{2}\pm\Re\sqrt{4m^{2}\eta^{2}\left(\widetilde{V}^{\mu\nu}S_{\nu}\right)^{2}-\mathcal{W}^{2}}$ (40)
holds true. Importantly, this subspace reduces to the EQED-type for massless axial-vector theories where $S^{2}=0$. The TRMW-type remains fully stable for Electroaxial theories involving massless pseudo-vectors, presenting intriguing and physically consistent characteristics for other massive axial-vector theories. Nonetheless, it permits the potential for chirality-asymmetric pair creation when a system with massive $S_{\mu}$ evolves into it from another subspace. Conversely, massive pseudo-vector theories within the NonAnom-type (including TRMW-type) may not be considered fundamental—primarily due to the presence of $\Pi_{\mu}S^{\mu}$ within $\mathcal{H}_{\mathrm{R}}$—and the condition $M_{S}\gg m$ must be satisfied along with constraints on $\mathcal{V}$ and/or $\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|$ to ensure the existence of at least one stable chirality sector. Further stability constraints may be required for different scenarios where $\mathrm{S}_{0}\left[A,S\right]$ could affect the consistency requirements or in cases where spacetime curvature is taken into account, or when both vector fields are quantized [64, *Nonpertrenorm, *StabilitydeRham].
IV The Exact One-Loop Effective Action
In this section, we derive closed-form expressions for effective actions of models situated within the previously classified consistent subspaces by employing zeta-function regularization. This process involves taking the trace in the eigenbasis of the Hessian and subsequently applying minimal subtraction (MS). We will initiate our analysis by categorizing the previously determined Hessians into two distinct classes and systematically identifying their eigenbasis on a case-by-case basis.
First class:
Since both fields are massless, we can utilize Fock-Schwinger’s gauge for each field, such that $V^{(R/L)}_{\mu}=-\left(eF_{\mu\nu}\pm\eta S_{\mu\nu}\right)x^{\nu}/2\pm\bar{S}_{\mu}$. If $F_{\mu\nu}$ and $S_{\mu\nu}$ form a regular pencil [*[See, e.g., ][forfurtherdetails.]gantmacher], it is possible to simultaneously transform them into Jordan’s form through an appropriate choice of coordinates. By applying boosts and rotations, we can choose a reference frame where each vector potential can be parameterized as
$V^{(H)}_{\mu}=\left(-E^{(H)}_{V}x,0,0,-B^{(H)}_{V}z,0\right)$ (41)
where $(H)=(R)$ or $(L)$ indicates the chirality sector. Here, the quantities $E^{(H)}_{V}$ and $B^{(H)}_{V}$ represent the electric and magnetic field analogs extracted from $V^{(H)}_{\mu\nu}=\partial_{\mu}V^{(H)}_{\nu}-\partial_{\nu}V^{(H)}_{\mu}$, for which the relevant invariants are closely related to Eq. (29),
$\begin{gathered}\mathcal{F}^{(H)}_{V}=\frac{1}{4}V^{(H)}_{\mu\nu}V_{(H)}^{\mu\nu}=e^{2}\mathcal{F}_{A}+\eta^{2}\mathcal{F}_{S}\pm\frac{e\eta}{2}S^{\mu\nu}F_{\mu\nu}\\ \mathcal{G}^{(H)}_{V}=-\frac{1}{4}\widetilde{V}^{(H)}_{\mu\nu}V_{(H)}^{\mu\nu}=e^{2}\mathcal{G}_{A}+\eta^{2}\mathcal{G}_{S}\pm\frac{e\eta}{2}\widetilde{F}^{\mu\nu}S_{\mu\nu}\end{gathered}$ (42)
and therefore
$\left(B^{(H)}_{V}\right)^{2}-\left(E^{(H)}_{V}\right)^{2}=2\mathcal{F}^{(H)}_{V}$ and $\mathbf{E}^{(H)}_{V}\mathbf{B}^{(R/L)}_{V}=\mathcal{G}_{V}^{(H)}$ (43)
Furthermore, since $\mathcal{V}=\mathcal{W}=0$, whereby we find $E^{(H)}_{V}=\pm B^{(H)}_{V}$ when $\mathcal{F}^{(H)}_{V}=\mathcal{G}^{(H)}_{V}=0$, that is, they form plane waves. Alternatively, if we allow the fields to take on complex values, we find $B^{(H)}_{V}=iE^{(H)}_{V}=\sqrt{\mathcal{F}^{(H)}_{V}}$. Thus, we can infer that Eq. (37) reduces to
$\mathcal{H}_{\mathrm{H}}=\left[p_{t}-E^{(H)}_{V}x\right]^{2}-p^{2}_{x}-\left[p_{y}-B^{(H)}_{V}z\right]^{2}-p_{z}^{2}-X_{(H)}$ (44)
where $X_{(R)}=-2\sqrt{\eta m\left|\widetilde{V}^{\mu\nu}S_{\mu}\right|}\,\sigma_{z}=\mathcal{X}_{+}\sigma_{z}$ and $X_{(L)}=-2i\sqrt{\eta m\left|\widetilde{V}^{\mu\nu}S_{\mu}\right|}\,\sigma_{z}=\mathcal{X}_{-}\sigma_{z}$. Next, by decomposing the $n$-th eigenstate of $\mathcal{H}$ into its chirality components $\left|\psi_{n}\right>=\left|\psi_{R,n}\right>+\left|\psi_{L,n}\right$>, we can express the trace as $\mathrm{Tr}e^{-i\tau\mathcal{H}}=\mathrm{Tr}_{\mathrm{R}}e^{-i\tau\mathcal{H}_{\mathrm{R}}}+\mathrm{Tr}_{\mathrm{L}}e^{-i\tau\mathcal{H}_{\mathrm{L}}}$ and employ momentum operators as
$\left[p_{t}-E^{(H)}_{V}x\right]^{2}=e^{\frac{ip_{t}p_{x}}{E^{(H)}_{V}}}\left[E^{(H)}_{V}x\right]^{2}e^{-\frac{ip_{t}p_{x}}{E^{(H)}_{V}}}$ (45)
allowing us to reduce Eq. (37) into quantum harmonic oscillators, such that
$\mathcal{H}_{\mathrm{H}}=-\left[p^{2}_{x}-\left(E^{(H)}_{V}\right)^{2}x^{2}\right]-\left[p^{2}_{z}+\left(B^{(H)}_{V}\right)^{2}z^{2}\right]-X_{(H)}$ (46)
Second class:
For more general theories, the Hessian can be expressed as
$\mathcal{H}=\Pi^{2}+\eta^{2}S^{2}+X$ (47)
where $X=\mathrm{diag}\left\{\mathcal{X}_{+}\sigma_{z},\mathcal{X}_{-}\sigma_{z}\right\}$, $X_{\pm}=\left|\lambda^{\left(\pm\right)}_{\pm}\right|$, and $\sigma_{z}=\mathrm{diag}\left\{1,-1\right\}$ represent the Pauli matrix in the $z$-direction. The next step involves decomposing $S_{\mu}$ into its transverse and longitudinal modes, followed by adopting an approximate Fock-Schwinger gauge for its transverse component within the domain defined by the quasi-static assumptions, specifically $S_{\mu}=S_{\mu\nu}x^{\nu}+\bar{S}_{\mu}$, satisfying $S^{2}=\left(S_{\mu\nu}x^{\nu}\right)^{2}+\bar{S}^{2}\approx\bar{S}^{2}$. We can extract $\bar{S}^{2}$ from the Hessian, leading us to define the effective mass as
$M^{2}=m^{2}-\eta^{2}\bar{S}^{2}$ (48)
This transformation reconfigures Eq. (47) into
$\mathcal{H}=\left(p_{\mu}-\frac{1}{2}eF^{\mu}_{\ \nu}x^{\nu}\right)^{2}+\eta^{2}\left(S^{\mu}_{\ \nu}x^{\nu}\right)^{2}+X$ (49)
This expression can be simplified under the conditions that either the matrix representations of both strength tensors commute and their Jordan forms exhibit non-zero blocks at the same positions, or that each tensor contains only one non-vanishing block. Conversely, if we parameterize the fields and apply the momentum operator, we obtain
$e^{-i\frac{p_{t}p_{x}}{a}}\left[\left(p_{t}-ax\right)^{2}+b^{2}t^{2}\right]e^{i\frac{p_{t}p_{x}}{a}}=a^{2}x^{2}+b^{2}\left(t+\frac{p_{x}}{a}\right)^{2}$
which essentially reinstitutes the original form, rather than yielding the desired result. Hence, we assume that the strength tensors possess appropriate Jordan blocks and define the analogous electric and magnetic fields derived from the transverse axial vector. We subsequently document the genuine electric and magnetic fields, both of which adhere to relations analogous to those in Eq. (43), specifically
$B^{2}_{A/S}-E^{2}_{A/S}=2\mathcal{F}_{A/S}$ and $\mathbf{E}_{A/S}\cdot\mathbf{B}_{A/S}=\mathcal{G}_{A/S}$ (50)
Consequently, it is feasible to parameterize $A_{\mu}=\left(-E_{A}x,0,-B_{A}z,0\right)$ and $S_{\mu}^{\perp}=\left(-E_{S}x,0,-B_{S}z,0\right)$ or interchangeably shift $z\to x$ for either vector. We can then apply the momentum operators as to shift
$x\to x+\frac{eE_{A}p_{t}}{E^{2}}$ $\text{ and }z\to z+\frac{eB_{A}p_{y}}{B^{2}}$ (51)
where we define $E^{2}=e^{2}E^{2}_{A}+\eta^{2}E^{2}_{S}$ and $B^{2}=e^{2}B^{2}_{A}+\eta^{2}B^{2}_{S}$. This transformation renders Eq. (49) into the following form:
$\mathcal{H}=\frac{\eta^{2}E^{2}_{S}}{E^{2}}p^{2}_{t}-\frac{\eta^{2}B^{2}_{S}}{B^{2}}p^{2}_{y}-p^{2}_{x}+E^{2}x^{2}-p^{2}_{z}-B^{2}z^{2}+X$ (52)
which is reminiscent of a Hamiltonian featuring two seemingly “massive” free directions, each characterized by different “mass” parameters, alongside the oscillatory components.
Now we can address both Hessian classes in a unified framework. Each class will exhibit Landau Levels [68] as eigenstates, with energy eigenvalues of the form
[…]
Here, $\lambda=\pm\frac{1}{2}$ corresponds to the spin eigenvalue, while $\mathcal{X}_{(R/L)}\leftrightarrow\mathcal{X}_{\pm}$. The index ($I=V,A,S$ or no index) represents the vector from which the “electromagnetic” fields have been derived. At this juncture, we only need to perform a standard textbook calculation [*[Ascanbeseeninsection33.4from][.]schwartz].
To facilitate our analysis, we assume integration within a Euclidean box of size $L$ and calculate the number of eigenstates contained within: (i) for the first class, each chirality sector may possess a different density of eigenstates, approximately given by $\sim 4\pi^{2}/\left(\mathcal{G}^{(H)}_{V}L^{2}\right)$, whereas (ii) for the second class, at energies below the fermion mass, the factors multiplying the momenta in Eq. (52) will have a minimal impact on the dispersion relation, leading to a density of eigenstates approximately $\sim 4\pi^{2}/\left(\mathcal{G}L^{2}\right)$, where $\mathcal{G}=EB$. Consequently, we derive
[…]
which, upon returning to the expression $L^{4}\to\int d^{4}x$ and Wick rotating $\tau\to i\tau$, ultimately yields
$\overline{\Gamma}_{c}\left[A,S\right]=\frac{1}{16\pi^{2}}\sum_{(H)}\int d^{4}x\ \frac{d\tau}{\tau}\left[\frac{e^{-M_{c}^{2}\tau}\mathcal{G}^{(H)}_{I}\cos{\left(\tau\mathcal{X}_{H}\right)}}{\sin{\left(\tau E^{(H)}_{I}\right)}\sinh{\left(\tau B^{(H)}_{I}\right)}}-\frac{e^{-m^{2}\tau}}{\tau^{2}}\right]$ (55)
Here, we denote the mass term of class $c$ as $M_{c}$, specifically $M_{1}=m$ or $M_{2}=M$. More explicitly, in the first class, the whole quantum correction vanishes for plane wave solutions, while it takes the form
[…]
$\left.+\frac{\left(E^{(L)}_{V}\right)^{2}\cos{\left(2\tau\sqrt{\eta m\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|}\right)}}{\sin^{2}{\left(\tau E^{(L)}_{V}\right)}}-\frac{2}{\tau^{2}}\right]$
for complexified fields, and results in
$\overline{\Gamma}_{2}\left[A,S\right]=\frac{1}{8\pi^{2}}\int d^{4}x\ \frac{d\tau}{\tau}\left[e^{-M^{2}\tau}\mathcal{G}\frac{\Re\cosh{\left(\tau\mathcal{X}_{+}\right)}}{\Im\cosh{\left(\tau\bar{\mathcal{X}}_{+}\right)}}-\frac{e^{-m^{2}\tau}}{\tau^{2}}\right]$ (57)
for the second class. For this, we have defined $\bar{\mathcal{X}}_{\pm}=\sqrt{2\mathcal{F}\pm 2i\mathcal{G}}=B\pm iE$, where $\mathcal{F}=\frac{1}{2}(B^{2}-E^{2})$. To employ a minimal subtraction scheme, we initially expand both components around $\tau\to 0$, resulting in
$\hskip-10.0pt\begin{aligned} \frac{\left(E^{(H)}_{I}\right)^{2}\cosh{\left(\tau\mathcal{X}_{+}\right)}}{\sin^{2}{\left(\tau E^{(H)}_{V}\right)}}&=\frac{1}{\tau^{2}}+\frac{1}{6}\left[2\left(E^{(H)}_{V}\right)^{2}\pm 12\eta m\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|\right]+\\ &+\frac{\tau^{2}}{120}\left[8\left(E^{(H)}_{V}\right)^{4}\pm 80\eta m\left(E^{(H)}_{V}\right)^{2}\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|+80\eta^{2}m^{2}\left(\widetilde{V}^{\mu\nu}S_{\nu}\right)^{2}\right]+\dots\end{aligned}$ (58)
for each chirality sector of the first class, and
$\mathcal{G}\frac{\Re\cosh{\left(\tau\mathcal{X}_{+}\right)}}{\Im\cosh{\left(\tau\bar{\mathcal{X}}_{+}\right)}}$ $=\frac{1}{\tau^{2}}+\frac{1}{12}\left[3\left(\mathcal{X}^{2}_{+}+\mathcal{X}^{2}_{-}\right)-4\mathcal{F}\right]+\frac{1}{720}\left[56\mathcal{F}^{2}-8\mathcal{G}^{2}-60\mathcal{F}\left(\mathcal{X}^{2}_{+}+\mathcal{X}^{2}_{-}\right)+\right$. (59)
$\left.+15\left(\mathcal{X}^{2}_{+}+\mathcal{X}^{2}_{-}\right)^{2}-30\mathcal{X}^{2}_{+}\mathcal{X}^{2}_{-}\right]\tau^{2}+\cdots$
for the second. By subtracting the first terms, we arrive at the generalized EH Lagrangians
[…]
$\left.+\frac{\left(E^{(L)}_{V}\right)^{2}\cos{\left(2\tau\sqrt{\eta m\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|}\right)}}{\sin^{2}{\left(\tau E^{(L)}_{V}\right)}}-\frac{2}{\tau^{2}}-\frac{1}{3}\left[\left(E^{(R)}_{V}\right)^{2}+\left(E^{(L)}_{V}\right)^{2}\right]\right\}$
and
$\mathcal{L}_{2}\left[A,S\right]=\frac{1}{8\pi^{2}}\int\frac{d\tau}{\tau}e^{-M^{2}\tau}\left[\mathcal{G}\frac{\Re\cosh{\left(\tau\mathcal{X}_{+}\right)}}{\Im\cosh{\left(\tau\bar{\mathcal{X}}_{+}\right)}}-\frac{1}{\tau^{2}}-\frac{1}{12}\left(3\mathcal{X}^{2}_{+}+3\mathcal{X}^{2}_{-}-4\mathcal{F}\right)\right]$ (61)
Before elaborating on this result, it is important to recognize that due to $S_{\mu}$ being a pseudo-vector, the mixed term in $\mathcal{F}_{V}^{(H)}$ ($\mathcal{G}_{V}^{(H)}$) does (not) violate CP invariance, contrasting with the corresponding term found in $\mathcal{V}$ ($\mathcal{W}$). Thus, additional constraints, such as $S^{\mu\nu}F_{\mu\nu}=\tilde{S}^{\mu\nu}F_{\mu\nu}=0$, may be necessary for model consistency [69]. If this constraint holds, we deduce that $\mathcal{F}^{(R)}_{V}=\mathcal{F}^{(L)}_{V}=\mathcal{V}$ and $\mathcal{G}^{(R)}_{V}=\mathcal{G}^{(L)}_{V}=\mathcal{W}$, which leads to the conclusion that Eq.(60) vanishes since it occurs when $\mathcal{V}=\mathcal{W}=0$, implying there are no surviving one-loop quantum corrections to the classical action. In other words, an effective action resembling the EH form for the TRMW-type subspace is exclusively feasible within CP-violating theories.
For the remainder of this section, we will operate within a MT-type subspace to elucidate these results, assuming $F_{\mu\nu}S^{\mu\nu}=\tilde{S}^{\mu\nu}F_{\mu\nu}=\left(\tilde{V}^{\mu\nu}S_{\nu}\right)^{2}=0$ for simplicity, where $X_{+}=\sqrt{2\mathcal{V}-4m^{2}\eta^{2}S^{2}+i\mathcal{W}}$. Thus, Eq. (61) transforms into
$\mathcal{L}_{MT}\left[A,S\right]=\frac{1}{8\pi^{2}}\int\frac{d\tau}{\tau}e^{-M^{2}\tau}\left[\mathcal{G}\frac{\Re\cosh{\left(\tau\mathcal{X}_{+}\right)}}{\Im\cosh{\left(\tau\bar{\mathcal{X}}_{+}\right)}}-\frac{1}{\tau^{2}}-\frac{2}{3}\left(e^{2}\mathcal{F}_{A}+\eta^{2}\mathcal{F}_{S}-2\eta^{2}m^{2}S^{2}\right)\right]$ (62)
and the fourth-order term of the expansion becomes
$\mathrm{I}_{4}=-\frac{4\tau^{2}}{45}\left[\left(e^{2}\mathcal{F}_{A}+\eta^{2}\mathcal{F}_{S}\right)^{2}+\frac{7}{4}\left(e^{2}\mathcal{G}_{A}+\eta^{2}\mathcal{G}_{S}\right)^{2}\right]+\frac{2\tau^{2}}{3}\eta^{4}m^{4}\bar{S}^{4}_{\mu}$ (63)
which aligns with the established EH Lagrangian for QED [2, 3, 7] in the limit as $\eta\to 0$. Additionally, we can verify this through perturbative analysis [40], expanding Eq. (7) to obtain
[…]
and maintaining careful consideration of the role of $\gamma^{5}$ during the application of dimensional regularization [*[Asfarasweknow, thisisanopenproblemconcerningdimensionalregularization.Nevertheless, weopttofollowtheprescriptionby][.]thooftveltman, *Collins]. The second-order term is given by
$\overline{\Gamma}^{(2)}\left[A,S\right]$ $=\frac{1}{\left(4\pi\right)^{2}}\int d^{4}xd^{4}y\frac{d^{4}p}{\left(2\pi\right)^{4}}e^{ip.\left(x-y\right)}\left[I_{1}\left(p^{2}g_{\mu\nu}-p_{\mu}p_{\nu}\right)V^{\mu}_{R}(y)V^{\nu}_{R}(x)+\right$. (65)
$\left.+m^{2}I_{2}\left(V^{\mu}_{R}(y)V^{L}_{\mu}(x)-V^{\mu}_{R}(y)V^{R}_{\mu}(x)\right)\right]+\begin{pmatrix}V_{R}\leftrightarrow V_{L}\\ \mathrm{P}_{R}\leftrightarrow\mathrm{P}_{L}\end{pmatrix}$
where
${\rm I}_{1}=\frac{\Delta}{6}-\int_{0}^{1}du(1-u)u\ln{\left[1-u(1-u)\frac{p^{2}}{m^{2}}\right]},$ ${\rm I}_{2}=\Delta-\int_{0}^{1}du\ln{\left[1-u(1-u)\frac{p^{2}}{m^{2}}\right]}$ (66)
[…]
$\left.-\frac{1}{15m^{2}}\left(e^{2}F_{\mu\nu}\Box F^{\mu\nu}+\eta^{2}S_{\mu\nu}\Box S^{\mu\nu}+5\eta^{2}S_{\mu}\Box^{2}S^{\mu}\right)+\mathcal{O}\left(\frac{1}{m^{4}}\right)\right]$
Thus, we conclude that Eq. (67) is consistent with the minimal subtraction term of Eq. (62) in the limit of constant fields since $\Delta_{M}=\int d\tau/\tau e^{-M^{2}\tau}=\Delta-\ln{\left(1-\eta^{2}S^{2}/m^{2}\right)}\approx\Delta$. It is noteworthy that the third-order term also conforms to expectations, despite a naive assumption that it might not be present if the analysis is kept merely at the expansion of Eq. (62), as will be elaborated in the following section.
V Chiral Anomaly
In this section, we explore the chiral anomaly and establish a connection between the perturbative approach and the exact EH action. This will also lay the foundations to utilize the pair creation rates derived in the next section in a straightforward model of baryogenesis.
Following the analysis in Eq. (64), we observe that at the third order, two terms survive while the others cancel due to Furry’s theorem. These terms correspond to the anomalous diagrams depicted in Fig.1. Specifically, they represent the linearly divergent integrals
$\overline{\Gamma}^{(3)}\left[A,S\right]$ $=\frac{-i\eta}{3}\int d^{4}x\,d^{4}y\,d^{4}z\frac{d^{4}k_{1}}{\left(2\pi\right)^{4}}\frac{d^{4}k_{2}}{\left(2\pi\right)^{4}}\frac{d^{4}p}{\left(2\pi\right)^{4}}e^{ik_{1}\cdot x+ik_{2}\cdot y}\delta\left(p-k_{1}-k_{2}\right)\times$ (68)
$\times\left(3e^{2}A^{\mu}_{x}A^{\nu}_{y}S^{\tau}_{z}I_{\mu\nu\tau}+\eta^{2}S^{\mu}_{x}S^{\nu}_{y}S^{\tau}_{z}I^{(5)}_{\mu\nu\tau}\right)$
where we adopt
[…]
By applying a Taylor series expansion and performing separated translations for each integration block $I_{(\cdot),\mu\nu\tau}$ from the first diagram, one can derive that [*[Thisprocedureisexploredinmoredetailatthechapter4of][.]Dobadolivro].
$I_{\mu\nu\tau}\left(a,b\right)-I_{\mu\nu\tau}\left(0,0\right)=-\frac{i\varepsilon_{\lambda\tau\nu\mu}\left(a-b\right)^{\lambda}}{8\pi^{2}}$ (70)
Figure 1: Anomalous triangle diagrams corresponding to the (a)
[…]
(++ permutations) and (b)
[…]
processes.
Thereby, by imposing Ward’s identities for vector current conservation, i.e., $\partial_{\mu}j^{\mu}=0\implies k_{1}^{\mu}I_{\mu\nu\tau}=k^{\mu}_{2}I_{\mu\nu\tau}=0$, in conjunction with the classical axial current’s continuity equation for massive fermions, given by $\partial_{\mu}j^{\mu}_{5}=2imj_{5}\implies p^{\lambda}I_{\mu\nu\lambda}=2imI_{\mu\nu}-\Delta_{\mu\nu}$, where
[…]
which is finite, and
[…]
$=\frac{1}{2\pi^{2}}\varepsilon_{\alpha\nu\beta\mu}k^{\alpha}_{2}k^{\beta}_{1}$
we can derive the well-known Adler–Bell–Jackiw anomaly [72, *belljackiw] (or Carroll–Field–Jackiw-Torsion [74] if the axial vector derives from torsion):
$\partial_{\mu}\left(j_{5}^{\mu}\right)_{(a)}=2im\left(j_{5}\right)_{(a)}-\frac{e^{2}}{8\pi^{2}}\tilde{F}^{\mu\nu}F_{\mu\nu}$ (73)
Here, $\left(j_{5}^{\mu}\right)_{(a)}=\frac{i}{3}A^{\mu}A^{\nu}I_{\mu\nu}$. A similar procedure can be applied to the other surviving diagram; however, in this case, one must impose symmetry between vertices instead of vector-current conservation to obtain
$\partial_{\mu}\left(j_{5}^{\mu}\right)_{(b)}=2im\left(j_{5}\right)_{(b)}-\frac{\eta^{2}}{24\pi^{2}}\tilde{S}^{\mu\nu}S_{\mu\nu}$ (74)
This indicates that the anomaly cancellation conditions of the Standard Model are sufficient to guarantee cancellation in this context as well [75, *anomalycancelation]. This insight presents an intriguing analogy to the pion within the weak field approximation by determining decay rates for each process (e.g., for coherent splitting):
$\Gamma\left(S\to\gamma\gamma\right)\propto\frac{e^{4}m^{3}}{256\pi^{3}}$ and $\Gamma\left(S\to SS\right)\propto\frac{\eta^{4}m^{3}}{2304\pi^{3}}$ (75)
It can be observed that $\Gamma\left(S\to\gamma\gamma\right)/\Gamma\left(S\to SS\right)=9e^{4}/\eta^{4}$. Thus, if $\eta<\sqrt{3}e$, there is a preferential decay of the axial field into a pair of photons. Such dynamics may suggest a mechanism for Primordial Magnetogenesis [19, *magnetobouncing] in cosmological models derived from theories incorporating propagating torsion [78, *torsioninflation]. Furthermore, the Eqs. (73) and (74) yield a term
$\overline{\Gamma}^{(3)}\left[A,S\right]=-\frac{\eta}{4\pi^{2}}\int d^{4}x\left[e^{2}S_{\mu}A_{\nu}\tilde{F}^{\mu\nu}-\frac{\eta^{2}}{6}\int d^{4}y\left(\partial_{\alpha}S^{\alpha}\right)_{x}{\Box}^{-1}_{xy}\left(S_{\mu\nu}\tilde{S}^{\mu\nu}\right)_{y}\right]$ (76)
for the (perturbative) effective action, where we adopt a mixed representation featuring a non-local component, expressed as $\Box^{-1}_{xy}=\int d^{4}p/(2\pi)^{4}e^{-ip\cdot(x-y)}1/p^{2}$, along with a 4D Chern-Simons-like term to underscore the topological nature of the component containing the electromagnetic field, which is a gauge field.
To address the Generalized EH action, we first need to investigate the functional form of the chiral current, i.e., $j^{\mu}_{5}=\delta\overline{\Gamma}/\delta S_{\mu}$. Let us denote the effective action as
$\Gamma\left[A,S\right]=\frac{1}{2}\int_{0}^{\infty}\frac{d\tau}{\tau}\ e^{-m^{2}\tau}f\left[A,S\right]-\frac{1}{8\pi^{2}}R$ (77)
where $f=\mathrm{Tr}e^{-\tau\mathcal{H}}$ and $R=\Delta_{M}\int d^{4}x\left(e^{2}F^{2}_{\mu\nu}+\eta^{2}S^{2}_{\mu\nu}+8\eta^{2}m^{2}S^{2}_{\mu}\right)/6$ is the minimal subtraction term for a MT-type subspace. It follows that $j^{\mu}_{5}$ is given by
[…]
$\left.-\frac{1}{12M^{2}\pi^{2}}\left(e^{2}F^{2}_{\alpha\beta}+\eta^{2}S^{2}_{\alpha\beta}+8\eta^{2}m^{2}S^{2}_{\mu}\right)\right]+\frac{8\Delta_{M}}{3}\eta^{2}m^{2}S^{\mu}$
where we utilized $\delta\Delta_{M}/\delta S_{\mu}=2\eta^{2}S^{\mu}/M^{2}$ for the contribution arising from $R$. The divergence leads to
$\partial_{\mu}j^{\mu}_{5}=\frac{1}{2}\int\frac{d\tau}{\tau}\ e^{-M^{2}\tau}\partial_{\mu}\frac{\delta}{\delta S_{\mu}}f$ (79)
since $\Pi_{\mu}S^{\mu}=0$. Notably, $\delta f/\delta S_{\mu}=-\tau\mathrm{Tr}\left[\eta\left(\left\{\gamma^{5},\Pi^{\mu}\right\}+2\eta S^{\mu}-2m\gamma^{\mu}\gamma^{5}\right)e^{-\tau\mathcal{H}}\right]$, allowing us to show a direct consistency of the effective action obtained within this subspace:
$\left(\partial_{\mu}j^{\mu}_{5}\right)_{\text{MT}}=\eta m\partial_{\mu}\left<\,\overline{\!{\psi}}\gamma^{\mu}\gamma^{5}\psi\right$> (80)
[…]
$j^{\mu}_{5}=\mathcal{N}\left[\left(j^{\mu}_{5}\right)_{\text{MT}}-\frac{M^{2}_{S}}{24\pi^{2}}S^{\mu}\right]$ (81)
where we employed the Proca action in Eq. (3) to establish $\Delta_{M}\int d^{4}x\delta S^{2}_{\alpha\beta}/\delta S^{\mu}=2\left(M^{2}_{S}S^{\mu}-j_{5}^{\mu}\right)$, with the normalization factor given by $\mathcal{N}=24\pi^{2}/\left(24\pi^{2}-\Delta_{M}\right)$. Its divergence results in
$\partial_{\mu}j^{\mu}_{5}=\mathcal{N}\left\{\left(\partial_{\mu}j^{\mu}_{5}\right)_{\text{MT}}+\left(j^{\mu}_{5}+\mathcal{L}S^{\mu}\right)\partial_{\mu}S^{2}+2\left[\eta^{2}\left(\partial_{\mu}S^{\mu}\right)^{2}+S^{\mu}\partial_{\mu}\right]\mathcal{L}-\frac{M^{2}_{S}}{24\pi^{2}}\left(\partial_{\mu}S^{\mu}\right)\right\}$ (82)
We identify contributions akin to Nieh–Yan’s term [80] ($\sim S^{\mu}\partial_{\mu}S^{2}$ in flat contorted spacetime), along with other contributions to the anomaly that encompass intricate expressions containing derivatives of the invariants. Indeed, aside from $\left(\partial_{\mu}j^{\mu}_{5}\right)_{\text{MT}}$, we need not employ the closed form of the Lagrangian obtained for the quasi-static/constant invariants approximation. Equation (82) indicates what contributions to the anomaly can be anticipated even in broader contexts involving varying fields, affirming that the result remains valid. This observation suggests that achieving stability outside of an MT-type subspace may require further considerations.
But this is not the complete story. To link Eqs. (73) and (74) to Eq. (62), we utilize the classical continuity equation $\left(\partial_{\mu}j^{\mu}_{5}\right)_{\text{MT}}=2imj_{5}$, where the latter term is defined as
$j_{5}=\left<\,\overline{\!{\psi}}\gamma^{5}\psi\right>\sim\frac{im}{16\pi^{2}}\int_{0}^{\infty}d\tau\ e^{-\tau M^{2}}\frac{\mathcal{G}}{\Im\cosh{\left(\tau\overline{X}_{+}\right)}}\ \mathrm{tr}\left(\gamma^{5}e^{-\frac{\tau}{2}V_{\mu\nu}\sigma^{\mu\nu}}\right)$ (83)
In this context, we focus on the term pertinent to the previously derived anomaly contributions within an MT-type subspace, under the condition $S_{\mu\nu}F^{\mu\nu}=\tilde{S}^{\mu\nu}F_{\mu\nu}=0$. Consequently, we find that $\mathrm{tr}\left(\gamma^{5}e^{-\frac{\tau}{2}V_{\mu\nu}\sigma^{\mu\nu}}\right)=-4im\Im\cosh{\left(\tau\overline{X}_{+}\right)}$, leading to
$\left(\partial_{\mu}j^{\mu}_{5}\right)_{\text{MT}}=2imj_{5}=\frac{m^{2}}{2\pi^{2}M^{2}}\int d\tau e^{-\tau M^{2}}\mathcal{G}\approx\frac{1}{2\pi^{2}}\int_{0}^{\infty}d\tau e^{-\tau M^{2}}\mathcal{G}$ (84)
for the regime $\eta^{2}S^{2}\ll m^{2}$. Ultimately, both Eq. (73) and Eq. (74) manifest after properly regulating Eq. (84) (for example, through the Pauli-Villars method [81]), although the resulting anomaly remains independent of the specific regularization approach employed [82]. A direct comparison with Eq. (80) also confirms this result, as
[…]
clearly contains the term $\int_{0}^{\infty}d\tau\,\eta^{2}e^{-\tau M^{2}}\mathcal{G}_{S}/(3\pi^{2})$, thereby agreeing with Eq. (84) given that the anomaly is proportional to $(\mathcal{G}-\eta^{2}\mathcal{G}_{S}/3)/(2\pi^{2})=(e^{2}\tilde{F}^{\mu\nu}F_{\mu\nu}+\eta^{2}\tilde{S}^{\mu\nu}S_{\mu\nu}/3)/(8\pi^{2})$.
One noteworthy consequence of this anomaly is as follows: Let us define $j^{\mu}=\left<\,\overline{\!{\psi}}\gamma^{\mu}\psi\right>=\left<\,\overline{\!{\psi}}\gamma^{\mu}\left({\rm P}_{R}+{\rm P}_{L}\right)\psi\right>=j^{\mu}_{R}+j^{\mu}_{L}$ and $j^{\mu}_{5}=j^{\mu}_{R}-j^{\mu}_{L}$. Since $A_{\mu}$ is a gauge field, its associated anomaly is inherently tied to topology, specifically, it is proportional to the winding number, which is an integer,
$\int d^{4}x\partial_{\mu}\left(j^{\mu}_{5}\right)_{(a)}=\int d^{4}x\left[\partial_{\mu}\left(j^{\mu}_{R}\right)_{(a)}-\partial_{\mu}\left(j^{\mu}_{L}\right)_{(a)}\right]=N_{R}-N_{L}=N_{5}$ (86)
In contrast, $S_{\mu}$ is not necessarily a gauge field itself, so Eq. (86) for the $S\to SS$ decay process can assume any real value. However, the existence of a $S\to\gamma\gamma$ decay channel implies both fields share topological content through their simultaneous coupling to the fermionic field. Specifically, when the couplings satisfy the condition for preferred $S\to\gamma\gamma$ decay, the axial vector is constrained to configurations that respect the topological structure inherited from the electromagnetic field. While $S_{\mu}$ need not exhibit classical instanton solutions, its fluctuations must be compatible with the winding number structure of $A_{\mu}$, effectively forcing the transferred particle-number asymmetry to obey the gauge field’s topological quantization. Remarkably, this constraint is stronger in NonAnom-type subspaces, which obey $e^{2}\mathcal{G}_{A}=\eta^{2}\mathcal{G}_{S}$ when $S^{\mu\nu}F_{\mu\nu}=0$.
Moreover, this anomaly indicates a violation of the conservation of particle number associated with each chirality following an instanton tunneling event, i.e.,
$\begin{gathered}\int d^{4}x\partial_{\mu}j^{\mu}=\int^{\infty}_{-\infty}dt\int d^{3}x\left[\partial_{\mu}\left(j^{\mu}\right)_{R}+\partial_{\mu}\left(j^{\mu}\right)_{L}\right],\\ =\left[N_{R}(\infty)-N_{R}(-\infty)\right]+\left[N_{L}(\infty)-N_{L}\left(-\infty\right)\right]=0\implies\Delta N_{R}=-\Delta N_{L}\end{gathered}$ (87)
which is a crucial requirement for scenarios such as baryogenesis [83]. We will leverage this insight in the subsequent section to explore particle creation and develop a toy model of pseudo-vector-driven baryogenesis as potential cause for the observed baryonic asymmetry in the contemporary universe.
VI Phenomenological Applications
In this section, we will exploit the non-unitarity of the effective action to derive the pair production rates. We will then apply these results within a simplified model of baryogenesis as a proof-of-concept. Furthermore, we will discuss potential extensions and implications of our findings in the realms of condensed matter physics and other related fields.
VI.1 Pair Production Rates
Let us begin by defining the Vacuum Transition Probability as $P_{\text{vac}}=\left|\left<A,S\right|\mathcal{S}_{\text{eff}}\left|A,S\right>\right|^{2}$, where $\mathcal{S}_{\text{eff}}=\exp{\left(-iW\left[A,S\right]\right)}$ is the S-matrix constructed from the full effective action. This probability quantifies the likelihood of remaining within the same background field configuration following scattering processes. According to the framework established by Schwinger [3], we observe that the quantity
$P_{\text{creation}}\left[A,S\right]=1-P_{\text{vac}}=1-\exp{\left(-2\Im W[A,S]\right)}$ (88)
provides a measure of the probability of exiting this vacuum state, thereby enabling the creation of any number of particles, therefore moving outside the scope of the effective action. This phenomenon can be modeled as a Poisson process [84], where the pair creation rate per unit volume and time is given by $\rho\left[A,S\right]=-2\Im\mathcal{L}$.
Although a nonzero vacuum transition probability may always exist, it becomes significant primarily in field configurations where the poles of $\mathcal{L}$ are dominant. For both classes of effective actions identified, these poles are related to the generalized analog of the electric field in a manner similar to the original EH action. Specifically, they occur at $\tau_{k}=\pi k/E_{I}^{(H)}$.
In the context of first-class Lagrangians, assuming B=iEB=-iE, the particle creation rates associated with each sector can typically be expressed as follows:
[…]
$\left.-2\sqrt{\eta m\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|}\ln{\left[1-\sinh{\left(\frac{2\pi}{E^{(R)}_{V}}\sqrt{\eta m\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|}\right)}e^{-\frac{\pi m^{2}}{E_{V}^{(R)}}}\right]}\right\}$
and
[…]
$\left.-2i\sqrt{\eta m\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|}\ln{\left[1-\sin{\left(\frac{2\pi}{E^{(L)}_{V}}\sqrt{\eta m\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|}\right)}e^{-\frac{\pi m^{2}}{E_{V}^{(L)}}}\right]}\right\}$
It is important to note that the last term in our expression is a pure imaginary number. Consequently, if this term dominates, such as when $4\sqrt{\eta m\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|}=E_{V}^{(L)}$, the vacuum state of left-handed fermions is stabilized, resulting in the absence of particle creation for this chirality. In contrast, for right-handed fermions, we observe an inevitable pair production phenomenon induced by a strong generalized electromagnetic field, irrespective of the value of $\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|$. Remarkably, Eq. (89) also demonstrates that the pair production rate of right-handed fermions can be effectively enhanced by adjusting the parameter $\sqrt{\eta m\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|}$.
When $\left|\widetilde{V}^{\mu\nu}S_{\nu}\right|=0$, the particle creation rates reduce to:
$\rho_{1}^{(H)}\left[E^{(H)}_{V}\right$] $=\frac{E^{(H)}_{V}}{8\pi^{2}}\left[\frac{E_{V}^{(H)}}{\pi}\sum_{k=1}^{\infty}\frac{1}{k^{2}}e^{-\frac{\pi km^{2}}{E_{V}^{(H)}}}-m^{2}\ln{\left(1-e^{-\frac{\pi m^{2}}{E_{V}^{(H)}}}\right)}\right]$ (91)
Although both chirality sectors can generate particle pairs, the specific field strengths given by $E^{(R/L)}_{V}=eE_{A}\pm\eta E_{S}$ reveal that one sector will require weaker fields to initiate vacuum decay. As a result, particle creation in this sector begins at an earlier energy, resulting in a net production of fermions with a preferred chirality. However, it is noteworthy that in this case, the chirality-asymmetric instability is not as pronounced as observed in the previous scenario.
Conversely, for the second class, we generally derive
$\rho_{2}\left[E,B,S\right]=\frac{EB}{4\pi^{2}}\sum_{k=1}^{\infty}\frac{(-1)^{k}e^{-\frac{\pi kM^{2}}{E}}}{k\sinh{\left(\frac{\pi kB}{eE}\right)}}\Re\cos{\left[\pi k\sqrt{1+2i\frac{B}{E}-\frac{\left(B^{2}+4\eta^{2}m^{2}S^{2}\right)}{E^{2}}}\right]}$ (92)
for the condition $\eta S\leq m$. This expression simplifies to the result obtained by Maroto [26] in the regime where $B=E_{S}=0$ and $2\eta mS\leq E$, yielding
$\rho_{\text{M}}\left[E_{A},S\right]=\frac{e^{2}E_{A}^{2}}{4\pi^{3}}\sum_{k=1}^{\infty}\frac{(-1)^{k}}{k^{2}}\cos{\left(\pi k\sqrt{1-\frac{4\eta^{2}m^{2}S^{2}}{e^{2}E_{A}^{2}}}\right)}\ e^{-\frac{\pi kM^{2}}{eE_{A}}}$ (93)
as expected. Notably, in the limits where $B^{2}\ll E^{2}$ and $4\eta^{2}m^{2}S^{2}\ll E^{2}$, we can reformulate Eq. (92) as
$\rho_{2}\left[E,B,S\right]=\frac{EB}{4\pi^{2}}\sum_{k=1}^{\infty}\frac{(-1)^{k}}{k}\coth{\left(\pi k\frac{B}{E}\right)}\cos{\left\{\pi k\left[1-\frac{\left(B^{2}+4\eta^{2}m^{2}S^{2}\right)}{2E^{2}}\right]\right\}}e^{-\frac{\pi kM^{2}}{E}}$ (94)
This formulation opens avenues for exploring scenarios such as the condition where $E_{S}=B_{A}=0$, $E_{A}\gg B_{S}$, and $E_{A}\gg 2\eta mS$, among other configurations. The corresponding vacuum transition probabilities for these rates, when contrasted with Schwinger’s original rate for QED, can be illustrated in FIG.2. It is noteworthy that when present, the axial vector tends to further stabilize the vacuum. Additionally, we can infer that a dynamical axial vector can create particles on its own. However, this effect is invariably accompanied by the term $S^{2}$, unless the axial vector is massless, indicating that its pair creation rates cannot perfectly replicate those of QED if it possesses mass.
Figure 2: Vacuum transition probabilities for an adiabatic laser pulse profile as described by Alkofer et al. [85], with pulse duration $\tau_{\rm pulse}=80\,fs$, and spatial extent $\sim\lambda=0.15\,nm$, according to (a) standard QED [3], (b) Eq. (91) for complexified fields, (c) Eq. (93) with $\eta S\sim 1\times 10^{17}\,eV$, and (d) Eq. (94) with $\eta B_{S}\sim 1\times 10^{15}\,eV/m$. Note the pronounced vacuum stabilization when the axial vector is present, indicating that pseudo-vector backgrounds suppress Schwinger pair production.
VI.2 Baryogenesis and Cosmology
Particle creation, closely associated with the chiral anomaly [86], also plays a pivotal role in providing CP violations; thus, such theories align with Sakharov’s conditions for baryogenesis. To delve into the feasibility of this scenario, we propose a toy model as a proof-of-concept.
Consider the presence of an axial vector (for instance, the pseudo-trace of torsion) with mass $M_{S}\gg m$ in the primordial universe. Assuming thermal equilibrium, once the temperature of the universe falls below $T\sim M_{S}$, the excitation of its new modes begins to decrease as $\propto e^{-M_{S}/T}$. This change allows us to model it as oscillating classically, with an energy density given by $\mathcal{E}\approx M_{S}+3T/2$.
At this juncture, the energy stored in this pseudo-vector is sufficiently substantial to initiate a particle creation process. This process will continue until the strength of its electric field analogue drops below a critical value or until it “freezes.” Subsequently, the remaining energy will undergo dilution until the present day. The asymmetry in particle numbers for each chirality at the conclusion of this process translates into a baryon asymmetry that remains conserved until the values recently measured [39].
Let us assume an appropriate coarse-graining process that preserves the Markovian approximation [87], allowing the particle creation rates to remain approximately constant at each step. From Eq. (86), it follows that $n_{5,0}=\int_{t_{i}}^{t_{f}}\rho\left[A,S\right]dt=N_{5}/V_{\text{comoving}}$ represents the total number of particles generated per co-moving volume at zero temperature, where $t_{i}$ and $t_{f}$ denote the initial and final instants of the process. At this time, the universe was dominated by radiation [88, *ParticleDataGroup], leading to
$t=\frac{0.301M_{\text{Planck}}}{\sqrt{g_{\ast}}T^{2}}\implies dt\approx-\frac{0.602M_{\text{Planck}}}{\sqrt{g_{\ast}}T^{3}}dT$ (95)
where $M_{\text{Planck}}=1.22\times 10^{19}\,GeV$ and $g_{\ast}\approx 106.75$ is taken as a constant. Assuming the entire process occurs prior to the electroweak symmetry breaking or even before neutrinos acquire mass (alternatively, we could consider calculating the temperature correction for massless fermions as a first approximation, given that their masses are small compared to $M_{S}$), we follow the prescription outlined by Kim et al. [90]. This involves assuming maximum efficiency in particle creation—thus using QED’s rate as an approximation—yielding
$n_{5}=\frac{1}{(2\pi)^{3}}\int n_{5,0}\tanh{\left(\frac{\omega}{2T}\right)}d^{3}\mathbf{k}=-\frac{-0.301\zeta\left(3\right)\eta^{2}E^{2}_{S}M_{\text{Planck}}}{\left(4\pi\right)^{4}\sqrt{g_{\ast}}}\Delta T$ (96)
where $\Delta T=T_{f}-T_{i}$ represents the temperature difference between the onset and conclusion of the process.
We can then leverage the fact that $n_{5}/s=\text{constant}$ to extrapolate the measured baryon-to-photon ratio $\eta_{B}=n_{5}/n_{\gamma}\approx 6.12\times 10^{-10}$, where $s\approx 7.04n_{\gamma}$ today, but was $s\approx(2\pi^{2}/45)g_{\ast}T_{f}^{3}$ during the radiation-dominated epoch. This allows us to find $n_{5}/s\approx 8.6932\times 10^{-11}\sim 10^{-10}$ in the past, leading to $n_{5}\approx(2\pi^{2}g_{\ast}T_{f}^{3}/45)\left(8.6932\times 10^{-11}\right)$. Furthermore, we assume that energy storage oscillates between the invariants of $S_{\mu}$, sharing it evenly on average as follows: $M^{2}_{S}\left<S^{2}\right>/2=-E^{2}_{S}/2=B^{2}_{S}/2=\mathcal{E}/3$. Thus,
$\eta\sqrt{\left<S^{2}\right>}=\frac{2\left(2\pi\right)^{3}T_{f}}{M_{S}\sqrt{45M_{\text{Planck}}}}\left(\frac{T_{i}}{T_{f}}-1\right)^{-\frac{1}{2}}\left(\frac{4g_{\ast}^{3}}{\zeta^{2}\left(3\right)}\right)^{\frac{1}{4}}\left(1.6974\times 10^{-5}\right)\,GeV^{\frac{3}{2}}$ (97)
Reorganizing in terms of scale as $T=\alpha M_{S}$, we derive:
$\eta\left<{S}\right>_{\text{rms}}\approx 1.54\frac{\alpha_{f}^{3/2}}{\sqrt{\alpha_{i}-\alpha_{f}}}10^{-11}\,GeV$ (98)
This implies that if the process concludes at temperatures in between $T_{f}\sim 10^{-2}M_{S}$ to $10^{-3}M_{S}$, and the temperature variation spans an order of $10^{-1}$ to $10^{1}$, then values ranging from $\eta\sqrt{\left<S^{2}\right>}=\eta\left<{S}\right>_{\text{rms}}\sim 10^{-14}\,GeV$ to $10^{-12}\,GeV$ would suffice to account for the present baryon asymmetry. Notably, these conditions yield expected values smaller than the total mass of neutrinos, thereby demonstrating consistency with the requirement that $m^{2}-\eta^{2}\left<S^{2}\right>\geq 0$ for well-behaved pair creation rates.
The density of these relics will dilute as $\propto a^{-3}$, where $a(T)\propto\left[T^{3}g_{\ast s}(T)\right]^{-1/3}$ is the scale factor. Thus, to obtain the relic density, we can perform
$\frac{\left<\mathcal{E}_{\text{relic}}\right>}{\left<\mathcal{E}(T_{f})\right>}=\frac{g_{\ast s}(T_{\text{today}})}{g_{\ast s}(T_{f})}\left[\frac{T_{\text{today}}}{T_{f}}\right]^{3}=\frac{\left<S^{2}_{\text{relic}}\right>}{\left<S^{2}\right>_{T_{f}}}$ (99)
given $T_{\text{today}}=2.73\,K=2.3\times 10^{-13}\,GeV$ and $g_{\ast s}(T_{\text{today}})=3.91$. We conclude that
$\eta\left<S_{\text{relic}}\right>_{\text{rms}}\lessapprox\frac{2.363}{M_{S}\sqrt{T_{i}-T_{f}}}\times 10^{-32}\,GeV^{\frac{5}{2}}=\frac{6.318}{m_{S}\sqrt{m_{S}\left(\alpha_{i}-\alpha_{f}\right)}}\times 10^{-61}\,GeV$ (100)
where we denote the mass in Planck units at the conclusion, i.e., $M_{S}=m_{s}M_{\text{Planck}}$. Ultimately, from $M^{2}_{S}\left<S^{2}\right>=2\mathcal{E}/3$, the relic energy density is approximately
$\mathcal{E}_{\text{relic}}\lessapprox\frac{6.8676}{\eta^{2}m_{S}\left(\alpha_{i}-\alpha_{f}\right)}10^{-83}\,GeV^{4}$ (101)
This relic energy density is remarkably low, comfortably situated well below current observational limits on Lorentz-violating interactions from astrophysical sources (e.g., $\gamma$-ray polarimetry) and laboratory tests [91, *Kosteleck_2011]. The corresponding value of the root-mean-square axial field strength $\eta\left<S_{\text{relic}}\right>_{\text{rms}}\lesssim 10^{-61}\,{\rm GeV}$ is also consistent with precision tests of CPT invariance. This parametric suppression arises naturally from the dilution factor $\sim(T_{\text{today}}/T_{f})^{3}$ combined with the stipulation that baryogenesis occurs before the electroweak symmetry breaking. In summary, despite the simplicity of this model, more rigorous constructions of such scenarios could indeed be viable explorations for future research.
It is noteworthy that the relic energy can be finely tuned to approximately $10^{-47}\,GeV^{4}$, while maintaining $\eta\left<S\right>_{\text{rms}}$ below the bounds imposed by experimental evidence. This observation indicates that such axial vector may also serve as a viable candidate for Dark Energy, given that massive vector fields can exhibit negative pressure [93]. In fact, the subset of parameter space where $\Pi_{\mu}S^{\mu}\neq 0$ could present intriguing possibilities for such theoretical proposals [94, *DarkProcaConstraints]. Nonetheless, it is essential to conduct a careful stability analysis to evaluate the implications of this hypothesis rigorously.
VI.3 Weyl Semimetals and Beyond
The electroaxial effective action derived above has direct applications to time-reversal asymmetric Weyl and Dirac semimetals, where axial-vector backgrounds emerge from Weyl-node separation $b_{\mu}=(b_{0},\mathbf{b})$ in momentum space, strain-induced pseudomagnetic fields, and magnetic textures [31, *Diractorsion2, *Diractorsion, 96, *torsionweylmetal, *EMWeylsemimetals]. In this context, the node separation vector plays the role of $S_{\mu}$, and our nonperturbative results become experimentally accessible: the required magnetic field strengths ($\sim 1$1010 T) are readily achievable in laboratory settings. Lorentz invariance is emergent rather than fundamental in these systems [35], permitting nonzero mixed terms $S_{\mu\nu}F^{\mu\nu}=\mathbf{E}_{A}\cdot\mathbf{B}_{S}+\mathbf{E}_{S}\cdot\mathbf{B}_{A}\neq 0$ that induce apparent vector-current nonconservation, requiring Bardeen-Zumino polynomial counterterms for consistent anomaly structure.
Our framework yields several testable predictions. First, the complete chiral-current divergence [Eqs. (82)]—which includes contributions beyond the standard Adler–Bell–Jackiw / Carroll–Field–Jackiw-Torsion [Eqs. (73),(74)] and Nieh–Yan terms—modifies anomalous transport coefficients and should manifest as corrections to longitudinal magnetoconductance in materials such as TaAs or Cd3As2. Second, the pair-production rates derived in Eqs. (91)–(94) translate into chiral-charge pumping rates between Weyl nodes under strong applied fields, with the vacuum-stabilization effect (Fig. 2) predicting enhanced threshold fields compared to naive Schwinger estimates, this effect should be observable in pulsed magnetic field experiments. Third, the electromagnetic-field generation via axial-vector decay [Eq. (75)] suggests that dynamical strain or magnetic textures in Weyl semimetals can source photon production, analogous to phenomena observed in magnetized plasmas [98, *plasma_axial2]. Quantitative comparison with transport measurements and pump-probe spectroscopy will be pursued in subsequent work.
Beyond condensed matter, two speculative extensions merit mention, though they require significant theoretical development. First, the longitudinal mode of a massive pseudo-vector $S_{\mu}$ could function as an axion-like coupling if the anomaly structure [Eq. (76)] is localized via auxiliary Stückelberg or Higgs fields, addressing the strong-CP problem as in axion models [100, 101, 102, 103]; our sector classification (Sec. III) would then constrain viable parameter space. Second, in Standard Model extensions with magnetic monopoles, the anomalous $S\to\gamma\gamma$ decay channel (Sec. V) could catalyze monopole-antimonopole pair production [20] in heavy-ion collisions, with potential observational implications for the MoEDAL [104] and RHIC [105] experiments; however, realizing this scenario requires promoting $A_{\mu}$ to a non-Abelian gauge field, placing it outside the scope of the present Abelian treatment.
VII Discussion and Outlook
In this work, we have presented a complete one-loop analysis of the electroaxial theory—Dirac fermions coupled to both electromagnetic and massive axial-vector backgrounds—within a controlled quasi-static approximation. By performing a full diagonalization of the functional Hessian, we classified the parameter space into physically distinct sectors, derived exact closed-form expressions for the EH effective action in the viable regimes, and extracted novel nonperturbative predictions for vacuum stability, pair production, and chiral anomaly structure. Our results unify and extend prior treatments revealing a robust vacuum-stabilization mechanism.
Nonetheless, several limitations inherent to our approach warrant acknowledgment. First, the current analysis lacks a thorough examination of the physical motivations underlying the necessity for a system to remain confined to one of the identified subspaces, as well as the reasons for its adherence to the stability criteria within that region of phase space. These constraints are likely application-dependent; however, for fundamental theories that incorporate axial vectors, there may not be an inherent justification for such restrictions. This potential absence of a priori reasoning could pose significant challenges for the applicability of the theoretical framework presented. Nevertheless, the framework offered here can be applied whenever a system visits the regions that we identify.
Second, our analysis assumes constant field invariants, which restricts applicability to homogeneous, slowly varying backgrounds. While standard for exact one-loop effective actions, this quasi-static approximation precludes direct modeling of rapidly time-dependent phenomena—such as ultrafast laser pulses or early-universe phase transitions—where non-adiabatic effects dominate. In such dynamical settings, one would expect contributions from topological terms like the Nieh–Yan invariant, which vanish identically in our constant-background limit but may become significant in curved or time-varying geometries.
Third, while we have verified consistency with perturbative expansions up to third order and matched known QED and Maroto limits, a full Functional Renormalization Group [106] analysis of the Electroaxial theoretic model at used (including higher-loop corrections and wave-function renormalization) remains open. This is particularly relevant for massive axial fields motivated by torsion [65], where consistency with quantum gravity or UV-completion scenarios (e.g., string theory) may impose additional constraints on couplings and mass scales.
Finally, our treatment is confined to flat spacetime. Gravitational effects, which are crucial for cosmological applications like baryogenesis or primordial magnetogenesis, are thus incorporated only indirectly via the axial field’s origin (e.g., as torsion’s pseudo-trace). A fully covariant generalization, incorporating curved backgrounds [107, *HeatKernelQG2] and spin-connection couplings, would be essential to assess backreaction, gravitational particle production [109], and the interplay between spacetime curvature and axial-induced anomalies [110].
Despite these limitations, our framework opens several promising avenues for future research in Condensed Matter Physics, Extensions of the Standard Model, Gravitation, and Cosmology. The exact results derived here not only resolve long-standing questions about stability and consistency but also furnish a toolkit for probing quantum vacuum structure from tabletop Weyl materials to the primordial universe. Therefore, we expect experimental verification of interesting new phenomena related to this work within the next years.
Appendix A Detailed Expressions of the Auxiliary Terms
In this appendix, we present the full expression of the terms from the characteristic equation and general solution presented in Sec. III. Let us begin with the terms from Eq. (27), which reads
$\begin{cases}\mathcal{A}=2\eta^{2}\left[m^{2}S^{2}-\left(\Pi_{\mu}S^{\mu}\right)^{2}\right]-\mathcal{V}\\ \mathcal{B}=\eta\mathcal{W}\Pi_{\mu}S^{\mu}\\ \mathcal{C}=\mathcal{A}^{2}+\mathcal{W}^{2}-4\eta^{2}\left[2\left(\Pi_{\mu}S^{\mu}\right)^{2}\mathcal{V}+m^{2}\left(\widetilde{V}^{\mu\nu}S_{\nu}\right)^{2}\right]\end{cases}$ (102)
The expression of $D_{\pm}$ in Eq. (28) is
$\mathcal{D}_{\pm}=\sqrt[3]{\Delta}-2\left(1\pm 3\right)\mathcal{A}+\frac{4}{\sqrt[3]{\Delta}}\left[4\mathcal{A}^{2}+24\eta^{4}m^{2}S^{2}\left(\Pi_{\mu}S^{\mu}\right)^{2}+3\mathcal{W}^{2}\right]$ (103)
with discriminant defined as
$\Delta=2\mathcal{K}+i\sqrt{2\mathcal{K}+64\left\{4\left(\mathcal{A}^{2}-\eta^{2}\Pi_{\mu}S^{\mu}\mathcal{V}\right)+3\mathcal{W}^{2}+12\eta^{2}m^{2}\left[\eta^{2}\left(\Pi_{\mu}S^{\mu}\right)S^{2}-\left(\tilde{V}^{\mu\nu}S_{\nu}\right)^{2}\right]\right\}^{3}}$ (104)
where
$\mathcal{K}=4\mathcal{A}^{3}-9\mathcal{A}\left[\mathcal{C}+64m^{2}\eta^{4}\left(\Pi_{\mu}S^{\mu}\right)^{2}S^{2}\right]-108\mathcal{B}^{2}$ (105)
Acknowledgements.
I thank I. L. Shapiro, G. P. de Brito, J. A. Helayël-Neto, W. Cesar e Silva, G. M. C. da Rocha, and P. M. M. de Souza for the discussions and useful comments. This work was partially supported by the Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq).